# Computing Generalized Ranks of Persistence Modules via Unfolding to Zigzag Modules

### v3

@@ 2.3 Idea of unfolding to zigzag module
Figure 2 illustrates this idea. The module $\mathbb{M}$ is defined on a four point poset $P=\{A,B,C,D\}$ shown on left. Bases of the vector spaces are shown in open brackets ‘()()( )’ and linear maps in these bases are shown as matrices. The poset is straightened into a zigzag path $A\rightarrow D\rightarrow C\leftarrow D\leftarrow B$. One way to look at this straightening is to view $P$ as a directed graph and to traverse all its vertices and edges sequentially possibly with repetition. Starting from the vertex $A$ and moving to an adjacent node disregarding the direction while noting down the visited node and the directed edge produces the zigzag path. This process of unfolding a poset into a zigzag path is formalized in section 5.1. The module $\mathbb{M}$ is unfolded into the zigzag module $\mathbb{M}_{ZZ}$ by copying the vector spaces and linear maps at vertices and edges respectively into their unfolded versions. For the module shown on the top, we get three interval modules (bars) in a decomposition of $\mathbb{M}_{ZZ}$, the full interval module supported on $[A,B]$ and the other two supported on two copies of $D$ respectively. Bases of one dimensional vector spaces for the interval modules are indicated beneath them. When we fold back $P_{ZZ}$ to $P$ (reversing the process of unfolding) sending $\mathbb{M}_{ZZ}$ to $\mathbb{M}$, the full interval module does fold back to a summand that is a full interval module because the vectors $v_{1}+v_{2}$111this is a vector addition, not a direct sum at two copies of $D$ are the same and hence map to the same vector in $\mathbb{M}(D)$. The other two interval modules supported on single-points in the decomposition of $\mathbb{M}_{ZZ}$ also fold back to a summand of $\mathbb{M}$ supported on a single point $D$ where $\mathbb{M}(D)$ is generated by the vector $v_{2}$.
@@ 3 Folding and Unfolding
For a folding $s:Q\rightarrow\mathrm{Fld}_{s}(Q)$ and a $Q$-module $\mathbb{N}$, we say $\mathbb{N}$ is $s$-foldable ((((or simply foldable)))) if $\mathbb{N}(q)=\mathbb{N}(q^{\prime})$ for every pair $q,q^{\prime}\in Q$ where $s(q)=s(q^{\prime})$.
@@
For a folding $s:Q\rightarrow\mathrm{Fld}_{s}(Q)$ and a $Q$-module $\mathbb{N}$, a summand $\mathbb{N}^{\prime}$ of $\mathbb{N}$ is called invertible if there exists a summand $\mathbb{Q}$ of $\mathbb{N}$ so that $\mathbb{N}^{\prime}\cong\mathbb{Q}$ and $\mathbb{Q}$ is foldable.
@@
Krull-Schmidt theorem of unique decomposition up to isomorphism [1] implies that $\overline{\mathbb{N}_{1}}\cong\mathbb{Q}$. Since $\mathbb{Q}$ is foldable, we have that $\overline{\mathbb{N}_{1}}$ is invertible.
Now, assume that $\overline{\mathbb{N}_{1}}$ is invertible. By definition, $\mathbb{N}=\mathbb{N}_{1}\oplus\mathbb{Q}$ for some foldable summand $\mathbb{Q}$ where $\overline{\mathbb{N}_{1}}\cong\mathbb{Q}$. Apply (1) above to conclude that $\mathrm{Fld}_{s}(\mathbb{N}_{1})$ exists and is a summand of $\mathbb{M}$.
@@ 4.1 Limit modules
We call a submodule $\mathbb{I}\subseteq\mathbb{M}$ a limit module for $\mathbb{M}:P\rightarrow\mathbf{vec}_{\mathbb{F}}$ if (i) $\mathbb{I}=\oplus_{i}\mathbb{I}^{I_{i}}$ where each $\mathbb{I}^{I_{i}}$ is an interval module with $I_{i}\cap I_{j}=\emptyset$ for $i\neq j$ and (ii) for every pair $p\leq_{P}q$, the morphism $\mathbb{I}(p\leq_{P}q)$ is either isomorphic to an identity or is a zero map. We call an ordered sequence of vectors $v_{p_{0}},\ldots,v_{p_{m}}$, $v_{p_{i}}\in\mathbb{M}(p_{i})$, a limit representative (or simply representative) if there is a limit module $\mathbb{I}$ for $\mathbb{M}$ so that $\mathbb{I}(p_{i})$ is spanned by the vector $v_{p_{i}}$ for each $i\in\{0,\ldots,m\}$. Conversely, every limit module $\mathbb{I}$ admits a representative ${\sf b}^{\mathbb{I}}_{p_{0}},\ldots,{\sf b}^{\mathbb{I}}_{p_{m}}$ where ${\sf b}^{\mathbb{I}}(p_{i})$ spans $\mathbb{I}(p_{i})$.
@@ 5 Zigzag module
A poset $P$ is called a zigzag poset iff there is a linear ordering $p_{0},\ldots,p_{m}$ of the points in $P$, called the zigzag path, so that for $i\in\{0,1,\ldots,m-1\}$, $p_{i}\leftrightarrow p_{i+1}$ are the only and all immediate pairs in $P$, i.e., zigzag path represents the Hasse diagram of $P$. We write $[p_{i},p_{j}]$ to denote an interval $I\subseteq P$ with the zigzag path $p_{i},p_{i+1},\cdots,p_{j}$.
@@
We will be interested in interval submodules $\mathbb{I}^{[b_{i},d_{i}]}$ of $\mathbb{M}_{ZZ}$. Such an interval module is either full or can be of four types determined by the types of its end points. The point $b_{i}$ is called closed if $b_{i}\not=0$ and the arrow between $b_{i-1},b_{i}$ is a forward arrow $\rightarrow$’ and is called open otherwise. Similarly, the point $d_{i}$ is called closed if $d_{i}\not=m$ and the arrow between $d_{i},d_{i+1}$ is a backward arrow and is called open otherwise. The interval module $\mathbb{I}^{[b_{i},d_{i}]}$ is called open-open, open-closed, closed-open, or closed-closed depending on whether $b_{i}$ and $d_{i}$ are both open, $b_{i}$ is open and $d_{i}$ closed, $b_{i}$ is closed and $d_{i}$ is open, or both $b_{i}$ and $d_{i}$ are closed respectively.
@@ 6 Algorithm
The algorithm (GenRank in pseudocode) takes a $P$-fitration $F$ and a degree $k$ for the homology group. First, it $s$-unfolds $P$ to a zigzag path of $P_{ZZ}$ and computes the filtration $F_{ZZ}$ (Step 1). Let $\mathbb{M}$ and $\mathbb{M}_{ZZ}$ be the modules obtained by applying the homology functor in degree $k$ on $F$ and $F_{ZZ}$ respectively as described above. We need to compute interval modules from $F_{ZZ}$ that represents a direct decomposition of $\mathbb{M}_{ZZ}$, i.e., we need a zigzag persistence algorithm that computes the intervals in the barcode with a representative (step 3). A sequence of $k$-cycles $z_{b},\ldots,z_{d}$ constitutes a representative of a limit module $\mathbb{I}:=\mathbb{I}^{[p_{b},p_{d}]}$ if $[z_{i}]$ is chosen as ${\sf b}^{\mathbb{I}}_{p_{i}}$ for $p_{i}\in[p_{b},p_{d}]$. The zigzag persistence algorithm in [15] can compute these representatives efficiently. Next, the algorithm checks how many interval modules in the computed decomposition of $\mathbb{M}_{ZZ}$ can be converted to $s$-complete modules which provides ${\sf{rk}}(\mathbb{M})$ according to the definition of $s$-completeness.
@@
    Let $\mathbb{I}$ be a convertible module where $\overline{\mathbb{I}}$ in $\mathcal{D}$ is invertible. Let $\mathcal{D}^{\prime}$ be the direct decomposition obtained from $\mathcal{D}$ by replacing $\mathbb{I}$ with the converted module (((($\mathrm{Convert}(\mathbb{I})$ in step 4 of GenRank)))), then $\kappa(\mathcal{D}^{\prime})=\kappa(\mathcal{D})+1$.
@@
For (2), observe that an interval module $\mathbb{J}\neq\mathbb{I}$ in $\mathcal{D}$ is foldable iff it is foldable in $\mathcal{D}^{\prime}$ because the only affected module in $\mathcal{D}$ is $\mathbb{I}$. We observe that $\overline{\mathbb{J}}$ remains invertible in $\mathcal{D}^{\prime}$ if and only if it were so in $\mathcal{D}$. This is because if there is a summand $\mathbb{Q}_{ZZ}$ of $\mathbb{M}_{ZZ}$ where $\mathbb{Q}_{ZZ}$ is foldable and $\overline{\mathbb{J}}\cong\mathbb{Q}_{ZZ}$ in $\mathcal{D}$, then $\overline{\mathbb{J}}$ remains isomorphic to $\mathbb{Q}_{ZZ}$ in $\mathcal{D}^{\prime}$ as well even if $\mathbb{I}$ changes to $\mathrm{Convert}(\mathbb{I})$. It follows that, compared to $\mathcal{D}$, the decomposition $\mathcal{D}^{\prime}$ has exactly one more foldable module, namely $\mathrm{Convert}(\mathbb{I})$, with its complement being invertible. Hence $\kappa(\mathcal{D}^{\prime})=\kappa(\mathcal{D})+1$. ∎
The algorithm GenRank draws upon Theorem 6.1. To accommodate computations with a fixed finite precision, we assume that the field $\mathbb{F}$ is finite. The algorithm takes every full interval module $\mathbb{I}$ that is not foldable and checks if it is convertible. If it is convertible, it converts it and then check if its complement $\overline{\mathbb{I}}$ for invertibility. It continues converting such modules to foldable modules until it cannot find any to convert. The current direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}$ changes with these conversions. At the end of this process, all convertible modules in the current decomposition whose complements are invertible become foldable themselves. Their number then coincides with ${\sf{rk}}(\mathbb{M})$. To determine the existence of the limit modules whose addition makes $\mathbb{I}$ foldable, we take the help of an annotation matrix $A_{p}$  [17, Chaper 4] for each complex $K_{p}$, $p\in P_{ZZ}$, computed in step 2. Without further elaborations, we only mention that annotation for a $k$-cycle $z\in Z_{k}(K_{p})$ (cycle group in degree $k$) is the coordinate of its class $[z]$ in a chosen basis of $H_{k}(K_{p})$; see [7] (annotation algorithm in [7] is described for $\mathbb{F}=\mathbb{Z}_{2}$ though it easily extends for any finite field $\mathbb{F}$). The representative cycles maintained for every interval module in the decomposition of $\mathbb{M}_{ZZ}$ at point $p$ form a cycle basis $[z_{1}],\ldots,[z_{g}]$ of $H_{k}(K_{p})$. We will see later that testing a full module for foldability amounts to calculating the annotations of the representative cycles of limit modules for certain points $p\in P_{ZZ}$ and performing certain matrix reductions with them; see section 6.1.
@@ 6.1 Convertibility of $\mathbb{I}$ and computing $\mathrm{Convert}(\mathbb{I})$
Assume that $\mathbb{I}$ is a full interval module in the current direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}=\bigoplus_{i}\mathbb{I}_{i}=\bigoplus_{i}\mathbb{I}^{[p_{b_{i}},p_{d_{i}}]}$. For each interval module $\mathbb{I}_{i}=\mathbb{I}^{[p_{b_{i}},p_{d_{i}}]}$ the algorithm computes a sequence of representative $k$-cycles $\{z^{i}_{p_{j}}\in Z^{k}(K_{p_{j}})\mid p_{j}\in\{p_{b_{i}},p_{b_{i}+1},\ldots,p_{d_{i}}\}\subseteq P_{ZZ}\}$. A representative for a limit module $\mathbb{I}^{[p_{b_{i}},p_{d_{i}}]}$ is given by the homology classes $[z^{i}_{p_{b_{i}}}],\ldots,[z^{i}_{p_{d_{i}}}]$. Such a representative for each limit module could be computed by an adaptation of the zigzag persistence algorithm in [30] or in [17, Chapter 4,section 4.3]. Instead, we apply the more efficient algorithm given in [15]. Then, a conversion of a convertible module $\mathbb{I}$ updates this representative as we update $\mathbb{I}$ to $\mathrm{Convert}(\mathbb{I})$.
@@ 6.2 Invertibility of $\overline{\mathbb{I}}$
If $\mathcal{D}$ is extendable, we have $\mathbb{M}_{ZZ}=\mathbb{I}\oplus\mathbb{Q}_{ZZ}$ by definition of extendability and the fact that $\mathcal{D}$ is a direct decomposition of $\mathbb{Q}_{ZZ}$. It follows that $\overline{\mathbb{I}}\cong\mathbb{Q}_{ZZ}$. Then, $\overline{\mathbb{I}}$ is invertible becasue $\mathbb{Q}_{ZZ}$ is foldable by construction.
@@
To prove the claim, observe that, due to the quotienting, the vector generating $\mathbb{I}(p)$ cannot lie in the vector space $\mathbb{Q}_{ZZ}(p)$ for every $p\in P_{ZZ}$. Then, $\mathbb{M}_{ZZ}(p)=\mathbb{I}(p)\oplus\mathbb{Q}_{ZZ}(p)$ for every $p\in P_{ZZ}$. Also, for every summand $\mathbb{I}^{\prime}$ of $\mathbb{Q}_{ZZ}$ in $\mathcal{D}$, the morphism $\mathbb{I}^{\prime}(p\leq_{P_{ZZ}}q)$ is a restriction of $\mathbb{M}_{ZZ}(p\leq_{P_{ZZ}}q)$. Therefore, for every interval module $\mathbb{I}_{i}\in\mathcal{D}\cup\{\mathbb{I}\}$, every morphism $\mathbb{I}_{i}(p\leq_{P_{ZZ}}q)$ is a restriction of $\mathbb{M}_{ZZ}(p\leq_{P_{ZZ}}q)$ over all pairs $p\leq_{P_{Z}Z}q$ while $\mathbb{M}_{ZZ}(p)=\oplus_{i}\mathbb{I}_{i}(p)$. It follows that $\mathcal{D}^{\prime}$ is a direct decomposition of $\mathbb{M}_{ZZ}$. ∎
We test if $\mathcal{D}$ is extendable by checking if the linear maps connecting two vectors (cycle classes) $[z_{p}]$ and $[z_{q}]$ for all immediate pairs $p,q\in P_{ZZ}$ in every interval module in $\mathcal{D}$ are preserved in $\mathcal{D}^{\prime}$. This means that if $p\leq_{P_{ZZ}}q$, the cycles $z_{p}$ and $z_{q}$ which are homologous in the complex $\hat{K}_{q}$ remain homologous in the complex $K_{q}$. This can be easily determined by checking the annotations of the cycles $z_{p}$ and $z_{q}$ in $K_{q}$. We need to compute the annotations of representative cycles for the interval modules in $\mathcal{D}$ in the complex $K_{q}$. We have already seen that it costs $O(tn^{\omega})$ time in total. Thus, the extendability of $\mathcal{D}$ can be checked in $O(tn^{\omega})$ time.

### v4

@@ 2.3 Idea of unfolding to zigzag module
Figure 2 illustrates this idea. The module $\mathbb{M}$ is defined on a four point poset $P=\{A,B,C,D\}$ shown on left. Bases of the vector spaces are shown in open brackets ‘$()$’ and linear maps in these bases are shown as matrices. The poset is straightened into a zigzag path $A\rightarrow D\rightarrow C\leftarrow D\leftarrow B$. One way to look at this straightening is to view $P$ as a directed graph and to traverse all its vertices and edges sequentially possibly with repetition. Starting from the vertex $A$ and moving to an adjacent node disregarding the direction while noting down the visited node and the directed edge produces the zigzag path. This process of unfolding a poset into a zigzag path is formalized in section 5.1. The module $\mathbb{M}$ is unfolded into the zigzag module $\mathbb{M}_{ZZ}$ by copying the vector spaces and linear maps at vertices and edges respectively into their unfolded versions. For the module shown on the top, we get three interval modules (bars) in a decomposition of $\mathbb{M}_{ZZ}$, the full interval module supported on $[A,B]$ and the other two supported on two copies of $D$ respectively. Bases of one dimensional vector spaces for the interval modules are indicated beneath them. When we fold back $P_{ZZ}$ to $P$ (reversing the process of unfolding) sending $\mathbb{M}_{ZZ}$ to $\mathbb{M}$, the full interval module does fold back to a summand that is a full interval module because the vectors $v_{1}+v_{2}$111this is a vector addition, not a direct sum at two copies of $D$ are the same and hence map to the same vector in $\mathbb{M}(D)$. The other two interval modules supported on single-points in the decomposition of $\mathbb{M}_{ZZ}$ also fold back to a summand of $\mathbb{M}$ supported on a single point $D$ where $\mathbb{M}(D)$ is generated by the vector $v_{2}$.
@@ 3 Folding and Unfolding
For a folding $s:Q\rightarrow\mathrm{Fld}_{s}(Q)$ and a $Q$-module $\mathbb{N}$, we say $\mathbb{N}$ is $s$-foldable ((or simply foldable)) if $\mathbb{N}(q)=\mathbb{N}(q^{\prime})$ for every pair $q,q^{\prime}\in Q$ where $s(q)=s(q^{\prime})$.
@@
For a folding $s:Q\rightarrow\mathrm{Fld}_{s}(Q)$ and a $Q$-module $\mathbb{N}$, a summand $\mathbb{N}^{\prime}$ in a decomposition $\mathbb{N}=\mathbb{N}^{\prime}\oplus\mathbb{N}$$^{\prime\prime}$ is called invertible if there exists a foldable summand $\mathbb{Q}$ of $\mathbb{N}$ so that $\mathbb{N}=\mathbb{Q}\oplus\mathbb{N}^{\prime\prime}$.
@@
Since $\mathbb{Q}$ is foldable, we have that $\overline{\mathbb{N}_{1}}$ is invertible.
Now, assume that $\overline{\mathbb{N}_{1}}$ is invertible. By definition, $\mathbb{N}=\mathbb{N}_{1}\oplus\mathbb{Q}$ for some foldable summand $\mathbb{Q}$. Apply (1) above to conclude that $\mathrm{Fld}_{s}(\mathbb{N}_{1})$ exists and is a summand of $\mathbb{M}$.
@@ 4.1 Limit modules
We call a submodule $\mathbb{I}\subseteq\mathbb{M}$ a limit module for $\mathbb{M}:P\rightarrow\mathbf{vec}_{\mathbb{F}}$ if (i) $\mathbb{I}=\oplus_{i}\mathbb{I}^{I_{i}}$ where each $\mathbb{I}^{I_{i}}$ is an interval module with $I_{i}\cap I_{j}=\emptyset$ for $i\neq j$ and (ii) for every pair $p\leq_{P}q$, the morphism $\mathbb{I}(p\leq_{P}q)$ is a surjection (this allows some particular types of interval modules to be limit modules, see e.g. Eq.(7)). We call an ordered sequence of vectors $v_{p_{0}},\ldots,v_{p_{m}}$, $v_{p_{i}}\in\mathbb{M}(p_{i})$, a limit representative (or simply representative) if there is a limit module $\mathbb{I}$ for $\mathbb{M}$ so that $\mathbb{I}(p_{i})$ is spanned by the vector $v_{p_{i}}$ for each $i\in\{0,\ldots,m\}$. Conversely, every limit module $\mathbb{I}$ admits a representative ${\sf b}^{\mathbb{I}}_{p_{0}},\ldots,{\sf b}^{\mathbb{I}}_{p_{m}}$ where ${\sf b}^{\mathbb{I}}(p_{i})$ spans $\mathbb{I}(p_{i})$.
@@ 5 Zigzag module
A poset $P$ is called a zigzag poset iff there is a linear ordering $p_{0},\ldots,p_{m}$ of the points in $P$, called the zigzag path, so that for $i\in\{0,1,\ldots,m-1\}$, $p_{i}\leftrightarrow p_{i+1}$ are the only and all immediate pairs in $P$, i.e., zigzag path represents the Hasse diagram of $P$. We write $p_{i},p_{j}]$ to denote an interval $I\subseteq P$ with the zigzag path $p_{i},p_{i+1},\cdots,p_{j}$.
@@
We will be interested in interval submodules $\mathbb{I}^{[b_{i},d_{i}]}$ of $\mathbb{M}_{ZZ}$. Such an interval module is either full or can be of four types determined by the types of its end points. The point $b_{i}$ is called closed if $b_{i}\not=0$ and the arrow between $b_{i-1},b_{i}$ is a forward arrow $\rightarrow$’ and is called open otherwise. Similarly, the point $d_{i}$ is called closed if $d_{i}\not=m$ and the arrow between $d_{i},d_{i+1}$ is a backward arrow and is called open otherwise. The interval module $\mathbb{I}^{[b_{i},d_{i}]}$ is called open-open, open-closed, closed-open, or closed-closed depending on whether $b_{i}$ and $d_{i}$ are both open, $b_{i}$ is open and $d_{i}$ closed, $b_{i}$ is closed and $d_{i}$ is open, or both $b_{i}$ and $d_{i}$ are closed respectively. Notice that a full interval module is considered open-open by definition.
@@ 6 Algorithm
The algorithm (GenRank in pseudocode) takes a $P$-fitration $F$ and a degree $k$ for the homology group. First, it $s$-unfolds $P$ to a zigzag path of $P_{ZZ}$ and computes the filtration $F_{ZZ}$ (Step 1). Let $\mathbb{M}$ and $\mathbb{M}_{ZZ}$ be the modules obtained by applying the homology functor in degree $k$ on $F$ and $F_{ZZ}$ respectively as described above. We need to compute interval modules from $F_{ZZ}$ that represents a direct decomposition of $\mathbb{M}_{ZZ}$, i.e., we need a zigzag persistence algorithm that computes the intervals in the barcode with a representative (step 3). A sequence of $k$-cycles $z_{b},\ldots,z_{d}$ constitutes a representative of a limit module $\mathbb{I}:=\mathbb{I}^{[p_{b},p_{d}]}$ if $z_{i}]$ is chosen as ${\sf b}^{\mathbb{I}}_{p_{i}}$ for $p_{i}\in[p_{b},p_{d}]$. The zigzag persistence algorithm in [15] can compute these representatives efficiently. Next, the algorithm checks how many interval modules in the computed decomposition of $\mathbb{M}_{ZZ}$ can be converted to $s$-complete modules which provides ${\sf{rk}}(\mathbb{M})$ according to the definition of $s$-completeness.
@@
    Let $\mathbb{I}$ be a convertible module where $\overline{\mathbb{I}}$ in $\mathcal{D}$ is invertible. Let $\mathcal{D}^{\prime}$ be the direct decomposition obtained from $\mathcal{D}$ by replacing $\mathbb{I}$ with the converted module $($$\mathrm{Convert}(\mathbb{I})$ in step 4 of GenRank)), then $\kappa(\mathcal{D}^{\prime})=\kappa(\mathcal{D})+1$.
@@
For (2), observe that an interval module $\mathbb{J}\neq\mathbb{I}$ in $\mathcal{D}$ is foldable iff it is foldable in $\mathcal{D}^{\prime}$ because the only affected module in $\mathcal{D}$ is $\mathbb{I}$. We observe that $\overline{\mathbb{J}}$ remains invertible in $\mathcal{D}^{\prime}$ if and only if it were so in $\mathcal{D}$. It follows that, compared to $\mathcal{D}$, the decomposition $\mathcal{D}^{\prime}$ has exactly one more foldable module, namely $\mathrm{Convert}(\mathbb{I})$, with its complement being invertible. Hence $\kappa(\mathcal{D}^{\prime})=\kappa(\mathcal{D})+1$. ∎
The algorithm GenRank draws upon Theorem 6.1. To accommodate computations with a fixed finite precision, we assume that the field $\mathbb{F}$ is finite. The algorithm takes every full interval module $\mathbb{I}$ that is not foldable and checks if it is convertible. If it is convertible, it converts it and then check if its complement $\overline{\mathbb{I}}$ for invertibility. It continues converting such modules to foldable modules until it cannot find any to convert. The current direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}$ changes with these conversions. At the end of this process, all convertible modules in the current decomposition whose complements are invertible become foldable themselves. Their number then coincides with ${\sf{rk}}(\mathbb{M})$. To determine the existence of the limit modules whose addition makes $\mathbb{I}$ foldable, we take the help of an annotation matrix $A_{p}$  [17, Chaper 4] for each complex $K_{p}$, $p\in P_{ZZ}$, computed in step 2. Without further elaborations, we only mention that annotation for a $k$-cycle $z\in Z_{k}(K_{p})$ (cycle group in degree $k$) is the coordinate of its class $[z]$ in a chosen basis of $H_{k}(K_{p})$; see [7] (annotation algorithm in [7] is described for $\mathbb{F}=\mathbb{Z}_{2}$ though it easily extends for any finite field $\mathbb{F}$). The representative cycles maintained for every interval module in the decomposition of $\mathbb{M}_{ZZ}$ at point $p$ form a cycle basis $z_{1}],\ldots,[z_{g}]$ of $H_{k}(K_{p})$. We will see later that testing a full module for foldability amounts to calculating the annotations of the representative cycles of limit modules for certain points $p\in P_{ZZ}$ and performing certain matrix reductions with them; see section 6.1.
@@ 6.1 Convertibility of $\mathbb{I}$ and computing $\mathrm{Convert}(\mathbb{I})$
Assume that $\mathbb{I}$ is a full interval module in the current direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}=\bigoplus_{i}\mathbb{I}_{i}=\bigoplus_{i}\mathbb{I}^{[p_{b_{i}},p_{d_{i}}]}$. For each interval module $\mathbb{I}_{i}=\mathbb{I}^{[p_{b_{i}},p_{d_{i}}]}$ the algorithm computes a sequence of representative $k$-cycles $\{z^{i}_{p_{j}}\in Z^{k}(K_{p_{j}})\mid p_{j}\in\{p_{b_{i}},p_{b_{i}+1},\ldots,p_{d_{i}}\}\subseteq P_{ZZ}\}$. A representative for a limit module $\mathbb{I}^{[p_{b_{i}},p_{d_{i}}]}$ is given by the homology classes $z^{i}_{p_{b_{i}}}],\ldots,[z^{i}_{p_{d_{i}}}]$. Such a representative for each limit module could be computed by an adaptation of the zigzag persistence algorithm in [30] or in [17, Chapter 4,section 4.3]. Instead, we apply the more efficient algorithm given in [15]. Then, a conversion of a convertible module $\mathbb{I}$ updates this representative as we update $\mathbb{I}$ to $\mathrm{Convert}(\mathbb{I})$.
@@ 6.2 Invertibility of $\overline{\mathbb{I}}$
If $\mathcal{D}$ is extendable, we have $\mathbb{M}_{ZZ}=\mathbb{I}\oplus\mathbb{Q}_{ZZ}$ by definition of extendability and the fact that $\mathcal{D}$ is a direct decomposition of $\mathbb{Q}_{ZZ}$. It follows that $\overline{\mathbb{I}}$ is invertible becasue $\mathbb{Q}_{ZZ}$ is foldable by construction.
@@
To prove the claim, observe that, due to the quotienting, the vector generating $\mathbb{I}(p)$ cannot lie in the vector space $\mathbb{Q}_{ZZ}(p)$ for every $p\in P_{ZZ}$. Then, $\mathbb{M}_{ZZ}(p)=\mathbb{I}(p)\oplus\mathbb{Q}_{ZZ}(p)$ for every $p\in P_{ZZ}$. Also, for every summand $\mathbb{I}^{\prime}$ of $\mathbb{Q}_{ZZ}$ in $\mathcal{D}$, the morphism $\mathbb{I}^{\prime}(p\leq_{P_{ZZ}}q)$ is a restriction of $\mathbb{M}_{ZZ}(p\leq_{P_{ZZ}}q)$ because we already established that $\mathbb{Q}_{ZZ}(p\leq_{P_{ZZ}}q)$ is a restriction of $\mathbb{M}_{ZZ}(p\leq_{P_{ZZ}}q)$. Therefore, for every interval module $\mathbb{I}_{i}\in\mathcal{D}\cup\{\mathbb{I}\}$, every morphism $\mathbb{I}_{i}(p\leq_{P_{ZZ}}q)$ is a restriction of $\mathbb{M}_{ZZ}(p\leq_{P_{ZZ}}q)$ over all pairs $p\leq_{P_{Z}Z}q$ while $\mathbb{M}_{ZZ}(p)=\oplus_{i}\mathbb{I}_{i}(p)$. It follows that $\mathcal{D}^{\prime}$ is a direct decomposition of $\mathbb{M}_{ZZ}$. ∎
We test if $\mathcal{D}$ is extendable by checking if the linear maps connecting two vectors (cycle classes) $z_{p}]$ and $z_{q}]$ for all immediate pairs $p,q\in P_{ZZ}$ in every interval module in $\mathcal{D}$ are preserved in $\mathcal{D}^{\prime}$. This means that if $p\leq_{P_{ZZ}}q$, the cycles $z_{p}$ and $z_{q}$ which are homologous in the complex $\hat{K}_{q}$ remain homologous in the complex $K_{q}$. This can be easily determined by checking the annotations of the cycles $z_{p}$ and $z_{q}$ in $K_{q}$. We need to compute the annotations of representative cycles for the interval modules in $\mathcal{D}$ in the complex $K_{q}$. We have already seen that it costs $O(tn^{\omega})$ time in total. Thus, the extendability of $\mathcal{D}$ can be checked in $O(tn^{\omega})$ time.